OpenStudy (anonymous):

A game is set up such that you have a 2/5 chance of winning $450 and a 3/5 chance of losing $100. What is your expected gain? A-120 B-70 C-45 D-30 I ended up getting C. Can someone else work this problem out and see if they also get it?

OpenStudy (tkhunny):

Expected Gain? How much is charged to play the game? How did you get C? And is that +45 or -45?

OpenStudy (anonymous):

No more information is given.

jimthompson5910 (jim_thompson5910):

2/5 chance of winning you win $450 so this multiplies out to (2/5)*450 = 180 ---------------------- 3/5 chance of losing $100 this multiplies out to (3/5)*(-100) = -60 ----------------------- Now add the two results: 180 + (-60) = 120 So you expect to gain $120 (on average) for any given turn